OpenStudy (anonymous):

if you have a spinner 1-4 and you spin it,what is the event space for landing on a even number

OpenStudy (anonymous):

if you have a spinner 1-4 and you spin it,what is the event space for landing on a even number

OpenStudy (anonymous):

find the probability ^_^, we have the followinf set of numbers from 1-4 1 2 3 4 number of even numbers in the set = 2 number of odd numbers in the set = 2 so the probability for landing on an even number is = (number of even numbers)/(total numbers found in the list) = 2/4 = 1/2 Correct me if I'm wrong ^_^

OpenStudy (anonymous):

I think the event space is {2,4} The sample space is {1,2,3,4}

OpenStudy (anonymous):

hmmm, so what are you trying to say? :)

OpenStudy (anonymous):

It is not asking for the probability. I think this person is just beginning to learn about probability and is learning the concept of sample space and event space. I may be really wrong, but just my thoughts.

OpenStudy (anonymous):

you might have a point, I just computed the following from the way I have understood the question ^_^ , you may be right since I myself am not fully convinced of my answer either.

OpenStudy (anonymous):

probabilty and possible outcomes im learning

OpenStudy (anonymous):

so I guessed the logic of the Question right? lol?

OpenStudy (anonymous):

The probablity is 1/2 The sample space is {1,2,3,4} I personally have never heard of the event space, but after reading online I think it is the subset of the sample space of the event you are trying to get. So I would say event space is {2,4}

OpenStudy (anonymous):

maybe the question didn't ask about "that" event space you're talking about? maybe there's more than one? possibly? ^_^